Sebastian’s math test

The default math mode font is math italic. This should not be confused with ordinary text italic. \mathbf produces bold face roman letters. If you wish to have bold face math italic letters, and bold face Greek letters and mathematical symbols, use the \boldmath command before going into math mode. This changes the default math fonts to bold. Greek is available in upper and lower case: α,β,Γ,Δ...ω,Ω

normal: x = 2π ⇒ x ≃ 6.28

mathbf x = 2π ⇒ x ≃ 6.28
boldmath x = 2π ⇒ x ≃ 6.28

There is also a calligraphic font for upper case letters; these are produced by the \mathcal command:ABCDE

ϕ(t) = \frac{1}{\sqrt{2π}} ∫_{0}^{t}e^{−x^{2}/2}dx  (1)

∏_{j≥0} ( ∑_{k≥0} a_{jk}z^{k}) = ∑_{k≥0} z^{n} ( ∑_{{k_{0},k_{1},...≥0 \atop k_{0}+k_{1}+...=n}} a_{0}k_{0}a_{1k_{1}}...)  (2)

π(n) = ∑_{m=2}^{n} ⌊ ( ∑_{k=1}^{m−1}⌊(m/k)/⌈m/k⌉⌋)^{−1} ⌋  (3)

\{\underbrace{\overbrace{a,...,a}^{k a′s},\overbrace{b,...,b}^{l b′s}}_{k+1 elements}\}  (4)

W^{+} \begin{matrix}↗ μ^{+} + ν_{μ} \\ → π^{+} + π^{0} \\ → κ^{+} + π^{0} \\ ↘ e^{+} + ν_{e} \end{matrix}

\frac{±| \begin{matrix}x_{1}− x_{2} & y_{1}− y_{2} & z_{1}− z_{2} \\ l_{1} & m_{1} & n_{1} \\ l_{2} & m_{2} & n_{2} \end{matrix} |}{\sqrt{| \begin{matrix}l_{1} & m_{1} \\ l_{2} & m_{2} \end{matrix} |^{2}+ | \begin{matrix}m_{1} & n_{1} \\ n_{1} & l_{1} \end{matrix} |^{2}+ | \begin{matrix}m_{2} & n_{2} \\ n_{2} & l_{2} \end{matrix} |^{2}}}

acute=ágrave=àddot=ätilde=ãbar=ābreve=ăcheck=ǎhat=âvec=a⃗ dot=ȧ

z = \sqrt{x}+ \frac{y}{a + b} + α_{2}+ \frac{Γ}{Ω}

z = \sqrt{x}+ \frac{y}{a + b} + α_{2}+ \frac{Γ}{Ω}

z = \sqrt{x}+ \frac{y}{a + b} + α_{2}+ \frac{Γ}{Ω}